# Transient QPOs of Fermi-LAT blazars with Linearly Multiplicative Oscillations

### v1

@@ Abstract
We present a study on the detection and characterization of transient quasi-periodic oscillations (QPOs) in the -ray emission of blazars 4C +31.03, MG1 J123931+0443, and PKS 1622253. Using light curves derived from Fermi Large Area Telescope data, we investigate oscillatory patterns characterized by periodic multiplicative amplitudes that vary linearly over time. By segmenting the light curves into increasing and decreasing trends, we analyze each segment independently, allowing for precise measurements of both the periodicity and long-term variations. To interpret these QPOs, we explore various theoretical scenarios that could explain their origin and underlying physical mechanisms. The estimated periods for blazars MG1 J123931+0443 and PKS 1622253 are consistent with precessional dynamics expected in binary supermassive black hole systems, although our current results do not allow for a definitive conclusion.
@@
BL Lacertae objects: general – galaxies: active – galaxies: nuclei1 Introduction
@@
Blazars are among the most intriguing and powerful types of extragalactic sources. These highly energetic objects are characterized by relativistic jets that are closely aligned with our line of sight (e.g., Wiita, 2006), making them prominent sources of radiation across the electromagnetic spectrum, from radio waves to rays (e.g., Urry, 2011; Peñil et al., 2024b). They can be divided into two subclasses, BL Lacertae objects and Flat Spectrum Radio Quasars (FSRQs), depending on the presence or absence of features in their optical spectra (Urry & Padovani, 1995). One of the most significant aspects of blazars is their variability across multiple timescales. While this variability is often associated with stochastic phenomena (Ruan et al., 2012; Covino et al., 2018), some blazars exhibit quasi-periodic oscillations (QPOs) in their emissions (e.g., Zhou et al., 2018; Adhikari et al., 2024). QPOs are of particular interest because they suggest the presence of coherent, cyclic processes within the blazar system, which can provide clues about the physical mechanisms governing their cores and relativistic jets. Potential explanations for QPOs include variations in the accretion rate (e.g., Gracia et al., 2003; Fichet de Clairfontaine et al., 2022) or dynamical interactions involving a secondary supermassive black hole (SMBH) within the blazar’s core (e.g., Sandrinelli et al., 2016; Sobacchi et al., 2017). Both scenarios can induce sustained, periodic variations in the observed emission.
@@
Recently, Peñil et al. (2025c) and Rico et al. (2024) reported blazars exhibiting long-term trends in their -ray emission (an increase or decrease in flux over time). Additionally, Rico et al. (2024) found that some of these blazars also display periodic patterns superimposed on these trends. Two distinct types of oscillations associated with these long-term variations were observed: additive oscillations, where the amplitude of the flux oscillations remains constant and independent of changes in the long-term trend (Peñil et al., 2024a), and multiplicative oscillations, where the amplitude scales proportionally with variations in the trend. These findings suggest that periodic behavior in blazars can manifest in different ways, potentially providing insights into the physical mechanisms governing their variability.
@@
In this paper, we analyze the QPOs in the -ray light curves (LCs) of three blazars using data from the Fermi Large Area Telescope (LAT, Atwood et al., 2009). These QPOs are transient, as they persist for a finite duration, typically spanning only a few cycles over several years. Their limited duration, relative to the total observational time, suggests that they are not continuous features but rather episodic phenomena that appear and disappear over time. These QPOs exhibit a distinctive trend in their oscillatory behavior: the amplitude of the oscillations varies over time, showing a gradual increase followed by a decrease. Different theoretical models have been proposed to explain the presence of such amplitude trends. One possibility is the interaction of an SMBH binary system with the surrounding gas, which could lead to changes in the accretion rate, thus affecting the amplitude of the QPOs over time (Adhikari et al., 2024). Another scenario involves magnetic reconnection events triggered by instabilities in the jet (Giannios, 2013a). By studying these blazars, we aim to expand our understanding of QPOs with multiplicative oscillations and their possible underlying physical origins.
@@ 2 Source Selection
We employed a sample of sources exhibiting long-term variability in their -ray emission. In particular, we analyzed 1,492 variable jetted AGN from the 4FGL-DR2 catalog (Ballet et al., 2020). This sample was systematically examined for long-term trends in Peñil et al. (2025c). Additionally, a subsample of these -ray LCs was further investigated in a detailed periodicity study by (Rico et al., 2024) using a novel method described in 3.
@@
A systematic study was performed by applying a pipeline consisting of different processing stages (Peñil et al., 2020, 2025a). As a result, 998 objects with a period with no statistical significance (1.5) were selected. Moreover, 337 objects were selected in the first analysis stage performed in Peñil et al. (2025c) for showing some evidence of the presence of a long-term trend, suggesting potential underlying processes that required further investigation (Peñil et al., 2025c). A cross-check of both subsamples resulted in a new total of 93 coincident objects.
@@
These 93 objects are further analyzed with the Continuous Wavelet Transform (CWT Torrence & Compo, 1998), a method that decomposes signals in both the time and frequency domains simultaneously. By applying the CWT, we can identify the presence of potential QPOs and locate the specific temporal segments of the LC where these QPOs occur. As a result, the CWT analysis of the sources reveals potential QPOs in specific LC regions, identifying them as transient QPO candidates. Through this analysis, we identify eleven blazars exhibiting transient QPO signatures. Among these, three blazars exhibited transient QPOs potentially explainable by the curved jet model (e.g., Sarkar et al., 2021; Prince et al., 2023). Subsequent analysis confirmed that two of these blazars were consistent with this jet model (Penil et al., 2025). In this paper, we refine our selection further by focusing specifically on three blazars, 4C +31.03, MG1 J123931+0443, and PKS 1622253, which show oscillations combined with linear trends (Table 1). The analyzed segments of their LCs are presented in Figure 1.
@@
For this analysis, we use the 28-day binned LCs above an energy of 0.1 GeV reported in Peñil et al. (2025c) and Rico et al. (2024). The 28-day binning allows us to capture and analyze long-term variations in the -ray data while reducing the impact of short-term variability.
@@ 3.1 Methods for the Search of QPOs
Two techniques are employed to identify QPOs in our blazar sample: the Singular Spectrum Analysis (SSA, Greco et al., 2016; Nina Golyandina, 2020) and the Generalized Lomb-Scargle Periodogram (GLSP, Zechmeister & Kürster, 2009).
SSA is a non-parametric method that decomposes a time series into its principal components, isolating trends, oscillating components, and noise. This decomposition allows for the identification of quasi-periodic components that may be hidden in the original noisy data (Rico et al., 2024). In QPO searches, SSA is advantageous because it does not require assumptions about the underlying periodicity, making it robust for detecting QPOs of varying frequencies and amplitudes.
Next, we apply the GLSP to this oscillatory component to determine the period and its associated significance, following the methodology outlined by Rico et al. (2024). The GLSP extends the traditional Lomb-Scargle Periodogram (LSP, Lomb, 1976; Scargle, 1982), which is widely used for detecting periodic signals in red-noise-dominated and unevenly-spaced data. GLSP improves the classic method by incorporating the ability to handle non-sinusoidal periodicities and by weighting data points according to their uncertainties. However, we note that it can be affected by irregularities in the amplitudes of the oscillations, as shown by Peñil et al. (2025b). This two-step approach improves our ability to detect periodic signals in blazars by effectively separating the underlying oscillations from noise.
Finally, the uncertainties associated with the periods reported by the GLSP and SSA are estimated using the half width at half maximum of the corresponding peaks (e.g., Otero-Santos et al., 2023).
@@ 3.2 Test Statistics of the QPOs
QPO searches in time-series data are often limited by noise, which can mask real periodic signals. Red noise, with its erratic brightness variations and increased power at lower frequencies, is a common issue in blazar time series that can lead to false QPO detections (Vaughan et al., 2016). Additionally, Poisson noise caused by the random detection of photons adds a white noise component, particularly affecting blazar LCs in -rays. These combined noise sources make it challenging to reliably identify true QPOs in observational data.
To address the challenges posed by the noise, we simulate 150,000 artificial LCs following the approach of Emmanoulopoulos et al. (2013). This method creates synthetic LCs that mimic the power spectral density (PSD) and probability density function (PDF) of the original time series. The PSD is modeled by a power law, , where is the normalization, is the spectral index, is frequency, and represents Poisson noise. These parameters are estimated using Maximum Likelihood Estimation and Markov Chain Monte Carlo (MCMC) analysis111We utilize the Python package emcee.. On the other hand, the PDF is described by a log-normal distribution.
@@ 3.3 Results
Using the previous methodology, we estimate the QPOs in the selected segments of our blazar sample. For 4C +31.03, the GLSP analysis suggests a primary peak at 52560 days (2.7) and a secondary peak at 28837 days (0.9) for SSA, while for GLSP, the primary peak is hinted at 51688 days (0.7), however, it does not show any significant signature of the secondary peak at 28837 days (0.1). Both methods report a similar period of 520 days but show notable differences in significance. Additionally, only the SSA hints at the presence of the secondary period at 288 days, with, however, a rather low significance.
The higher confidence in SSA suggests it effectively reduces noise, enhancing the peak’s significance. We also note that, as reported by Peñil et al. (2025b), amplitude changes in recurrent or periodic flares such as the ones observed here can have a very strong impact on the significance derived for such QPO patterns.
@@ 4 Theoretical Models
QPOs observed in the -ray LCs of blazars have been extensively studied, as they provide valuable insights into the dynamic processes occurring in the cores of these objects. Recently, Peñil et al. (2025c) reported that some blazars exhibit an associated long-term trend in their emission. These trends influence the amplitude of both aperiodic and periodic oscillations by modulating their multiplicative values over time as the long-term trend evolves. In some cases, QPOs appear superimposed on these long-term variations, suggesting a complex interplay between periodic oscillations and secular emission changes (Rico et al., 2024). This type of QPO suggests that they may result from multiple interdependent processes, indicating a complex system where independent phenomena could interact to generate QPOs and modulate their amplitudes multiplicatively.
@@
The intrinsic temporary structure of the QPOs suggests that they could arise from a range of transient phenomena. For instance, accretion disk instabilities can lead to fluctuations in the mass accretion rate (Abdo et al., 2010), which, when propagating inward, modulate the energy injection into the jet. Fluctuations grow as they propagate through the disk and into the jet (e.g., Rieger, 2019), reaching peak modulation when they fully transfer energy to the jet base. If disk instabilities occur periodically, they can generate QPOs that naturally follow the observed amplitude evolution (Czerny, 2004).
@@
In another scenario, proposed by Wiita (2011), QPOs can arise from non-axisymmetric structures in the jet, most likely in the magnetic field. Because of this, when a shock wave moves outward through the jet, it interacts with these irregularities, causing variations in the emitted light. When the shock moves through a helical structure within the jet, the effect is similar to a change in the jet’s direction, resulting in an apparent shift in the viewing angle. If the jet is cylindrical and the shock velocity remains constant, the recurring boosts in flux would generate a periodic signal. In this case, the QPO persists if the shock traverses the helical structure and retains the same period if another similar event occurs. Additionally, Wiita (2011) suggested that recurring boosts in the turbulent flow behind a propagating shock can also produce QPOs. If a dominant turbulent cell is present in the jet, Doppler boosting should be observed, leading to a QPO. However, as the turbulent cell gradually decays, the QPO amplitude will diminish over time due to the intrinsic stochastic nature of turbulence. Furthermore, the period of the QPO may vary since the size and/or velocity of the turbulent cell can also change.
@@
Another scenario could involve magnetohydrodynamic (MHD) instabilities that can induce helical distortions in the jet, leading to periodic emission variations (Dong et al., 2020). The instability starts weakly, grows as the magnetic field lines twist further, reaches a peak when it is fully developed, and then declines as the field structure dissipates. MHD instabilities can produce magnetic reconnections, leading to the formation of plasma blobs within the jet, also known as plasmoids (Sironi et al., 2016). During reconnection events, the formation and interaction of plasmoids can lead to rapid and intense -ray flares. As these plasmoids merge or are ejected from the reconnection site, they can produce variability patterns in the jet emission (Shukla & Mannheim, 2020), associated with rapid (duration of hours to days) flare activities (Meyer et al., 2021) or variations on timescales of months to years (Christie et al., 2019). In addition to that, the plasmoids have also been proposed to be the origin of QPOs observed in the -ray emission of blazars. Specifically, in the model proposed by Mohan & Mangalam (2015), a plasmoid follows a helical path determined by the magnetic field geometry. Specifically, the plasmoid experiences a combination of toroidal and poloidal magnetic field components, leading to a spiral trajectory. The pitch angle of the helix, which governs the tightness of the spiral, is dependent on the relative strength of these magnetic components (Asada et al., 2008). Moreover, the size of the plasmoid plays a role in the emitted flux. Larger plasmoids have longer light-crossing times, leading to extended flare durations and more symmetric LCs (Finke, 2024). In contrast, smaller plasmoids produce shorter, sharper flares, which can result in more rapid variability in the LCs (Meyer et al., 2021). Additionally, the speed of the plasmoid within the jet significantly impacts the observed flux. Faster-moving plasmoids experience stronger Doppler boosting, leading to an enhancement in the observed intensity, whereas slower plasmoids contribute less prominently to the variability pattern (Meyer et al., 2021). Additionally, energy dissipation and flux decay are tied to plasmoid dynamics. Flux decay can be influenced by the radiative cooling of particles within the plasmoid and its expansion, which reduces its magnetic field strength and particle density, leading to a decrease in emission intensity over time (Christie et al., 2019).
@@
As the plasmoid moves through the jet, it may encounter discontinuities in physical properties, such as velocity shear or heterogeneous-density regions. These interactions can generate shock waves, which, in turn, compress the plasma and enhance particle acceleration (Giannios, 2013b). This process leads to bursts of synchrotron and inverse Compton emission (Marscher & Gear, 1985; Heavens & Drury, 1988). As a result, the amplitude of oscillations in the emission flux can increase due to the shocks amplifying the density and energy of pre-existing fluctuations (Fichet de Clairfontaine et al., 2022). The oscillation amplitude peaks when the shock achieves maximum efficiency in energizing particles, gradually declining as energy dissipates and the shock weakens. This energy dissipation could result in a progressive decrease in oscillation amplitude, marking the transition from an active acceleration phase to a more quiescent state (Fichet de Clairfontaine et al., 2022).

### v3

@@ Abstract
We present a study on the detection and characterization of transient quasi-periodic oscillations (QPOs) in the -ray emission of blazars 4C +31.03, MG1 J123931+0443, and PKS 1622253. Using light curves derived from Fermi Large Area Telescope data, we investigate oscillatory patterns characterized by periodic multiplicative amplitudes that vary linearly over time. By segmenting the light curves into increasing and decreasing trends, we analyze each segment independently, allowing for precise measurements of both the periodicity and long-term variations. To interpret these QPOs, we explore various theoretical scenarios that could explain their origin and underlying physical mechanisms. The variability observed in 4C +31.03 is more consistent with a stochastic process, whereas the periods estimated for MG1 J123931+0443 and PKS 1622253 align with the precessional dynamics expected from binary supermassive black hole systems. However, the current results remain tentative and do not allow for a definitive conclusion.
@@
Blazar – galaxies: active – galaxies: nuclei1 Introduction
@@
Blazars are among the most intriguing and powerful types of extragalactic sources. These highly energetic objects are characterized by relativistic jets that are closely aligned with our line of sight (e.g., wiita_lecture), making them prominent sources of radiation across the electromagnetic spectrum, from radio waves to rays (e.g., urry_multiwavelengh; penil_2020; penil_mwl). They can be divided into two subclasses, Flat Spectrum Radio Quasars (FSRQs) and BL Lacertae objects, depending on the presence or absence of features in their optical spectra (urry_n). One of the most significant aspects of blazars is their variability across multiple timescales. While this variability is often associated with stochastic phenomena (ruan_stocastic; covino_negation), some blazars exhibit quasi-periodic oscillations (QPOs) in their emissions (e.g., zhou_PKS_2247_131; sagar_pg1553). In practical terms, a QPO is typically identified when at least three cycles are observed with peak separations consistent (e.g., vaughan_criticism; penil_kink_2025). These criteria allow for performing a study of the observed modulation that could reflect a physically meaningful variability rather than random stochastic fluctuations.
QPOs are of particular interest because they suggest the presence of coherent, cyclic processes within the blazar system, which can provide clues about the physical mechanisms governing their cores and relativistic jets. Potential explanations for QPOs include variations in the accretion rate (e.g., gracias_modulation_disk; fichet_increase_trends_shock_2022) or dynamical interactions involving a secondary supermassive black hole (SMBH) within the blazar’s core (e.g., sandrinelli_redfit; sobacchi_binary). Both scenarios can induce sustained, periodic variations in the observed emission.
@@
Recently, penil_2025 and alba_ssa reported blazars exhibiting long-term trends in their -ray emission (an increase or decrease in flux over time). Additionally, alba_ssa found that some of these blazars also display periodic patterns superimposed on these trends. Two distinct types of oscillations associated with these long-term variations were observed: additive oscillations, where the amplitude of the flux oscillations remains constant and independent of changes in the long-term trend (penil_2022_multiwave), and multiplicative oscillations, where the amplitude scales proportionally with variations in the trend. These findings suggest that periodic behavior in blazars can manifest in different ways, potentially providing insights into the physical mechanisms governing their variability.
@@
In this paper, we analyze the QPOs in the -ray light curves (LCs) of three blazars using data from the Fermi Large Area Telescope (LAT, fermi_lat). These QPOs are transient, as they persist for a finite duration, typically spanning only a few cycles over several years. Their limited duration, relative to the total observational time, suggests that they are not continuous features but rather episodic phenomena that appear and disappear over time. These QPOs exhibit a distinctive trend in their oscillatory behavior: the amplitude of the oscillations varies over time, showing a gradual increase followed by a decrease. Different theoretical models have been proposed to explain the presence of such amplitude trends. One possibility is the interaction of an SMBH binary system with the surrounding gas, which could lead to changes in the accretion rate, thus affecting the amplitude of the QPOs over time (sagar_pg1553). Another scenario involves magnetic reconnection events triggered by instabilities in the jet (giannios_trend_magnetic). By studying these blazars, we aim to expand our understanding of QPOs with multiplicative oscillations and their possible underlying physical origins.
@@ 2 Source Selection
We employed a sample of sources exhibiting long-term variability in their -ray emission. In particular, we analyzed 1,492 variable jetted AGN from the 4FGL-DR2 catalog (4fgl_dr2). This sample was systematically examined for long-term trends in penil_2025. Additionally, a subsample of these -ray LCs was further investigated in a detailed periodicity study by (alba_ssa) using a novel method described in 3.
@@
A systematic study was performed by applying a pipeline consisting of different processing stages (penil_2025_24; penil_2025_4fgl). As a result, 998 objects with a period with no statistical significance (1.5) were selected (penil_2025_4fgl). Moreover, 337 objects were selected in the first analysis stage performed in penil_2025 for showing some evidence of the presence of a long-term trend111This evidence comes from performing periodic analyses on both detrended and non-detrended LCs, allowing us to identify frequency-domain changes potentially introduced by long-term trends (penil_2025)., suggesting potential underlying processes that required further investigation (penil_2025).
A cross-check of both subsamples resulted in a new total of 93 coincident objects. Although the periods and trends identified in penil_2025_4fgl; penil_2025 were statistically not significant, in this work, we revisit the previous subset as candidates of interest. These sources show potential indications of structured variability as periodic oscillations and long-term coherence that justify further examination beyond the initial automated classification.
@@
These 93 objects are further analyzed with the Continuous Wavelet Transform (CWT wavelet_torrence), a method that decomposes signals in both the time and frequency domains simultaneously. By applying the CWT, we can identify the presence of potential QPOs and locate the specific temporal segments of the LC where these QPOs occur. As a result, the CWT analysis of the sources reveals potential QPOs in specific LC regions, identifying them as transient QPO candidates. Through this analysis, we identify eleven blazars exhibiting transient QPO signatures. Among these, three blazars exhibited transient QPOs potentially explainable by the curved jet model (e.g., sarkar_curved_jet; anuvab_curve_jet). In this scenario, periodic variations arise from changes in the viewing angle, which modulate the observed flux through differential Doppler boosting as emitting regions travel along the curved geometry of the jet. Subsequent analysis confirmed that two of these blazars were consistent with this jet model (penil_curve_jet_2025). In this paper, we refine our selection further by focusing specifically on three FSRQs, 4C +31.03, MG1 J123931+0443, and PKS 1622253, which show oscillations combined with linear trends (Table 1). The analyzed segments of their LCs are presented in Figure 1. The remaining five sources exhibit more complex QPO structures that cannot be adequately described by the linear-trend framework used in this study. Their analysis will require exploring alternative scenarios, such as quadratic long-term trends (quadratic_trend_binary_combi), or exponential rise-decay oscillations (exponential_shape).
The remaining objects of the sample present a more complex QPO structure, which will require the analysis of the application of alternative scenarios as quadratic trends (quadratic_trend_binary_combi).
@@
For this analysis, we use the 28-day binned LCs above an energy of 0.1 GeV reported in penil_2025 and alba_ssa. The 28-day binning allows us to capture and analyze long-term variations in the -ray data while reducing the impact of short-term variability. Details of the Fermi-LAT data reduction procedure are provided in penil_2025_4fgl.
@@ 3.1 Methods for the Search of QPOs
Two techniques are employed to identify QPOs in our blazar sample: the Singular Spectrum Analysis (SSA, ssa_greco; SSA_algorithm) and the Generalized Lomb-Scargle Periodogram (GLSP, lomb_gen).
SSA is a non-parametric method that decomposes a time series into its principal components, isolating trends, oscillating components, and noise. This decomposition allows for the identification of quasi-periodic components that may be hidden in the original noisy data (alba_ssa). In QPO searches, SSA is advantageous because it does not require assumptions about the underlying periodicity, making it robust for detecting QPOs of varying frequencies and amplitudes.
Next, we apply the GLSP to this oscillatory component to determine the period and its associated significance, following the methodology outlined by alba_ssa. The GLSP extends the traditional Lomb-Scargle Periodogram (LSP, Lomb_1976; Scargle_1982), which is widely used for detecting periodic signals in red-noise-dominated and unevenly-spaced data. GLSP improves the classic method by incorporating the ability to handle non-sinusoidal periodicities and by weighting data points according to their uncertainties. However, we note that it can be affected by irregularities in the amplitudes of the oscillations, as shown by penil_flare_2025. This two-step approach improves our ability to detect periodic signals in blazars by effectively separating the underlying oscillations from noise.
Finally, the uncertainties associated with the periods reported by the GLSP and SSA are estimated using the half width at half maximum of the corresponding peaks (e.g., otero_mwl).
@@ 3.2 Test Statistics of the QPOs
QPO searches in time-series data are often limited by noise, which can mask real periodic signals. Red noise, with its erratic brightness variations and increased power at lower frequencies, is a common issue in blazar time series that can lead to false QPO detections (vaughan_criticism). Additionally, Poisson noise caused by the random detection of photons adds a white noise component, particularly affecting blazar LCs in -rays. These combined noise sources make it challenging to reliably identify true QPOs in observational data.
To address the challenges posed by the noise, we simulate 150,000 artificial LCs following the approach of emma_lc. This method creates synthetic LCs that mimic the power spectral density (PSD) and probability density function (PDF) of the original time series. The PSD is modeled by a power law, , where is the normalization, is the spectral index, is frequency, and represents Poisson noise. These parameters are estimated using Maximum Likelihood Estimation and Markov Chain Monte Carlo (MCMC) analysis222We utilize the Python package emcee.. On the other hand, the PDF is described by a log-normal distribution.
@@ 3.3 Results
Using the previous methodology, we estimate the QPOs in the selected segments of our blazar sample. For 4C +31.03, the GLSP analysis suggests a primary peak at 52560 days (2.7) and a secondary peak at 28837 days (0.9), while for SSA, the primary peak is hinted at 51688 days (0.7); however, it does not show any significant signature of the secondary peak at 28837 days (0.1). Both methods report a similar period of 520 days but show notable differences in significance. Additionally, only the SSA hints at the presence of the secondary period at 288 days, with, however, a rather low significance.
The higher confidence in SSA suggests it effectively reduces noise, enhancing the peak’s significance. We also note that, as reported by penil_flare_2025, amplitude changes in recurrent or periodic flares such as the ones observed here can have a very strong impact on the significance derived for such QPO patterns.
@@ 4 Theoretical Models
QPOs observed in the -ray LCs of blazars have been extensively studied, as they provide valuable insights into the dynamic processes occurring in the cores of these objects. Recently, penil_2025 reported that some blazars exhibit an associated long-term trend in their emission. These trends influence the amplitude of both aperiodic and periodic oscillations by modulating their multiplicative values over time as the long-term trend evolves. In some cases, QPOs appear superimposed on these long-term variations, suggesting a complex interplay between periodic oscillations and secular emission changes (alba_ssa). This type of QPO suggests that they may result from multiple interdependent processes, indicating a complex system where independent phenomena could interact to generate QPOs and modulate their amplitudes multiplicatively.
4.1 Accretion Disk-Origin Scenarios
@@
The intrinsic temporary structure of the QPOs suggests that they could arise from a range of transient phenomena. For instance, accretion disk instabilities can lead to fluctuations in the mass accretion rate (abdo_variability), which, when propagating inward, modulate the energy injection into the jet. Fluctuations grow as they propagate through the disk and into the jet (e.g., rieger_2019), reaching peak modulation when they fully transfer energy to the jet base. If disk instabilities occur periodically, they can generate QPOs that naturally follow the observed amplitude evolution (czerny_accretion_disk).
4.2 Jet-Origin Scenarios
@@
QPOS can also originate from processes intrinsic to the jet itself. In the scenario proposed by wiita_cells_2011, QPOs can arise from non-axisymmetric structures in the jet, most likely in the magnetic field. Because of this, when a shock wave moves outward through the jet, it interacts with these irregularities, causing variations in the emitted light. When the shock moves through a helical structure within the jet, the effect is similar to a change in the jet’s direction, resulting in an apparent shift in the viewing angle. If the jet is cylindrical and the shock velocity remains constant, the recurring boosts in flux would generate a periodic signal. In this case, the QPO persists if the shock traverses the helical structure and retains the same period if another similar event occurs. Additionally, wiita_cells_2011 suggested that recurring boosts in the turbulent flow behind a propagating shock can also produce QPOs. If a dominant turbulent cell is present in the jet, Doppler boosting should be observed, leading to a QPO. However, as the turbulent cell gradually decays, the QPO amplitude will diminish over time due to the intrinsic stochastic nature of turbulence. Furthermore, the period of the QPO may vary since the size and/or velocity of the turbulent cell can also change.
4.2.1 Magnetohydrodynamic Instabilities
@@
Another scenario could involve magnetohydrodynamic (MHD) instabilities that can induce helical distortions in the jet, leading to periodic emission variations (dong_kink_2020). The instability starts weakly, grows as the magnetic field lines twist further, reaches a peak when it is fully developed, and then declines as the field structure dissipates. MHD instabilities can produce magnetic reconnections, leading to the formation of plasma blobs within the jet, also known as plasmoids (plasmoids_sinori_2016). During reconnection events, the formation and interaction of plasmoids can lead to rapid and intense -ray flares. As these plasmoids merge or are ejected from the reconnection site, they can produce variability patterns in the jet emission (magnetic_reconnections_plasmoids_shukla_2020), associated with rapid (duration of hours to days) flare activities (meyer_sizeblobs_2021) or variations on timescales of months to years (timescale_plasmoids_christie_2019).
In addition to that, the plasmoids have also been proposed to be the origin of QPOs observed in the -ray emission of blazars. Specifically, in the model proposed by mohan_mangalam_blobs, a plasmoid follows a helical path determined by the magnetic field geometry. Specifically, the plasmoid experiences a combination of toroidal and poloidal magnetic field components, leading to a spiral trajectory. The pitch angle of the helix, which governs the tightness of the spiral, is dependent on the relative strength of these magnetic components (asada_pitch_angle). Moreover, the size of the plasmoid plays a role in the emitted flux. Larger plasmoids have longer light-crossing times, leading to extended flare durations and more symmetric LCs (finke_size_blobs_2024). In contrast, smaller plasmoids produce shorter, sharper flares, which can result in more rapid variability in the LCs (meyer_sizeblobs_2021). Additionally, the speed of the plasmoid within the jet significantly impacts the observed flux. Faster-moving plasmoids experience stronger Doppler boosting, leading to an enhancement in the observed intensity, whereas slower plasmoids contribute less prominently to the variability pattern (meyer_sizeblobs_2021). Additionally, energy dissipation and flux decay are tied to plasmoid dynamics. Flux decay can be influenced by the radiative cooling of particles within the plasmoid and its expansion, which reduces its magnetic field strength and particle density, leading to a decrease in emission intensity over time (timescale_plasmoids_christie_2019).
@@
As the plasmoid moves through the jet, it may encounter discontinuities in physical properties, such as velocity shear or heterogeneous-density regions. These interactions can generate shock waves, which, in turn, compress the plasma and enhance particle acceleration (giannios_blobs_envelope_2013). This process leads to bursts of synchrotron and inverse Compton emission (shock_marscher_1985; shock_heavens_1988). As a result, the amplitude of oscillations in the emission flux can increase due to the shocks amplifying the density and energy of pre-existing fluctuations (fichet_increase_trends_shock_2022). The oscillation amplitude peaks when the shock achieves maximum efficiency in energizing particles, gradually declining as energy dissipates and the shock weakens. This energy dissipation could result in a progressive decrease in oscillation amplitude, marking the transition from an active acceleration phase to a more quiescent state (fichet_increase_trends_shock_2022).
4.2.2 External Jet Agents